Find the unknown currents 14 in amperes for the network. Given: I1=20A; 15= 12A; 16= 8A. 11 a 15 b W с WW 16 I,
Answer #1

I=20A a Is = 12 A I-8A wmt G MW WH с -8A Iz Apply Kch at Node a Si=15+ Iz 20 17 SA Apply KCL at nodec Is = Ist Iz 12=8+ I Iz=